OpenStudy (givingtree):

REALLY NEED HELP FOR UNDERSTANDING. Dealing with an Adjustable Rate Mortgage and need to figure out monthly payments. Please help, I really need to know how to do this!

OpenStudy (givingtree):

Adjustable Rate Mortgage: 3% with terms 5/1 with a 2/6 cap for 30 years (Assume the interest rate increases by 1.25% after the initial period and every 10 years thereafter.) ARM loan of : $130,000

OpenStudy (givingtree):

The first # in the term 5/1 : is the initial interest rate period. The second # in the term 5/1 : tells how often the interest rate can adjust. So, the mortgage will have a fixed rate for 5 years, after which it can adjust each year. The 2/6 cap means the interest rate can adjust up to 2% for each adjustment period, but will never exceed 6% more than the initial interest rate. This means, the maximum interest rate for this mortgage is 9%. I feel this first part of the information is correct, but it needs to be checked please.

OpenStudy (givingtree):

@TheSmartOne ! Could you help me with this ARM?

OpenStudy (givingtree):

anyone?

OpenStudy (triciaal):

sorry would have to think

OpenStudy (givingtree):

sorry, this is a lot. just any help I can get, trying to finish up senior year.

OpenStudy (givingtree):

@imqwerty

OpenStudy (3mar):

It is not from honesty to answer a question I do not know. I really don't know. Sorry for disappointing you!

OpenStudy (givingtree):

no no! thank you for letting me know @3mar !!!

jimthompson5910 (jim_thompson5910):

you have the correct interpretation so far @givingtree

OpenStudy (givingtree):

Thank you. I just don't know how to set it up, like what number would i put in for the i variable, interest?

jimthompson5910 (jim_thompson5910):

most likely i = 0.03/12 = 0.0025 since I'm assuming the i refers to the monthly interest rate. It's not 100% clear though.

OpenStudy (givingtree):

i = periodic interest rate = \[\frac{ r }{ n }\]

jimthompson5910 (jim_thompson5910):

ok so r = 0.03 and n = 12

OpenStudy (givingtree):

it might be safer to do a fraction in this formula?

jimthompson5910 (jim_thompson5910):

decimal form may be easier

OpenStudy (givingtree):

true.

OpenStudy (givingtree):

let me set up the problem and check back with you to see if it's correct?

OpenStudy (givingtree):

I can pay you through paypal.

OpenStudy (givingtree):

?

jimthompson5910 (jim_thompson5910):

http://www.bankrate.com/calculators/mortgages/adjustable-rate-mortgage-arm-calculator.aspx this is telling me that there are 4 different payment values, so I'm not sure if your teacher wants all 4 or just the initial payment or maybe the max payment?

OpenStudy (givingtree):

Getting a picture for you.

OpenStudy (givingtree):

does that help at all?

jimthompson5910 (jim_thompson5910):

That table is what your teacher gave you?

OpenStudy (givingtree):

yes, that i put on paper.

jimthompson5910 (jim_thompson5910):

This piece of info `Adjustable Rate Mortgage: 3% with terms 5/1 with a 2/6 cap for 30 years` and this piece of info `(Assume the interest rate increases by 1.25% after the initial period and every 10 years thereafter.) ` don't match up fully

OpenStudy (givingtree):

I don't know what to say, it's just what i was given .-.

jimthompson5910 (jim_thompson5910):

but I'm assuming your teacher wants you to go with `(Assume the interest rate increases by 1.25% after the initial period and every 10 years thereafter.)`

jimthompson5910 (jim_thompson5910):

because the other option has it increasing every year

OpenStudy (givingtree):

So, what should I do?

jimthompson5910 (jim_thompson5910):

I'd go with the second one. Anyways, I'm assuming you know the fixed rate mortgage payment formula?

OpenStudy (givingtree):

absolutely

jimthompson5910 (jim_thompson5910):

what is the monthly payment for those first five years (year 1 through year 5) ?

OpenStudy (givingtree):

i haven't solved for that part yet, or any. um lemme put the formula together real quick to check

OpenStudy (givingtree):

what would be the i ?

OpenStudy (givingtree):

Like, will the value of i, (aka r/n) change for each period? 1-5 6-15 16-25 26-30

OpenStudy (givingtree):

or stay the same?

jimthompson5910 (jim_thompson5910):

for year1 through year5, i = 0.03/12 = 0.0025

OpenStudy (givingtree):

and after year1 through year5, i would change to?

jimthompson5910 (jim_thompson5910):

r would increase 1.25% each time year1 through year5: r = 3% -----> r = 0.03 year6 through year15: r = 4.25% -----> r = 0.0425 year16 through year25: r = 5.5% ----> r = 0.055 year26 through year30: r = 6.75% ----> r = 0.0675 i = r/n

jimthompson5910 (jim_thompson5910):

n = 12 the whole time

OpenStudy (givingtree):

Perfect thank you!!! I'll get to solving now!

OpenStudy (givingtree):

\[m = \frac{ (130000)\frac{ .03 }{ 12 }(1+\frac{ .03 }{ 12 } ) ^{360} }{ (1+\frac{ .03 }{ 12 } )^{360} -1 }\]

OpenStudy (givingtree):

did i set this up correctly?

jimthompson5910 (jim_thompson5910):

yes you did

OpenStudy (givingtree):

yay! i'll solve to find all the year periods now. brb

OpenStudy (givingtree):

hey, I really appreciate you for taking all this time to help me. almost done!!

jimthompson5910 (jim_thompson5910):

what is the monthly payment for the first 60 months?

OpenStudy (givingtree):

$548.09 for year1-year5

jimthompson5910 (jim_thompson5910):

good how much interest is paid for the first month? hint: monthly interest rate = (3%)/12 = 0.25%

OpenStudy (givingtree):

uummmmmmmm

OpenStudy (givingtree):

one sec *

OpenStudy (givingtree):

interest rate for a month is going to be 0.025?

jimthompson5910 (jim_thompson5910):

interest due for given month = (monthly interest rate in decimal form)*(balance owed) interest due for month1 = (0.0025)*(130,000) interest due for month1 = $325

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (givingtree):

the (.03/12) was just simply reduced to .0025, yes makes sense

jimthompson5910 (jim_thompson5910):

so $325 is the amount in interest to be paid $548.09 is the monthly payment 548.09-325 = 223.09 is the principal paid

OpenStudy (givingtree):

this stuff is so much...

jimthompson5910 (jim_thompson5910):

after paying off 223.09 in principal, the balance goes from $130,000 to $129,776.91 this is all shown in the attached image on the highlighted line

jimthompson5910 (jim_thompson5910):

it's a lot given there are 360 of these payments to make. Which is why calculators and spreadsheets are very handy

OpenStudy (givingtree):

i bet!!

jimthompson5910 (jim_thompson5910):

again I used http://www.bankrate.com/calculators/mortgages/adjustable-rate-mortgage-arm-calculator.aspx to make up that payment schedule

OpenStudy (givingtree):

bookmarking that thank you.

OpenStudy (givingtree):

so the monthly payment will not change?

jimthompson5910 (jim_thompson5910):

for month1 through month60, the payment is 548.09

jimthompson5910 (jim_thompson5910):

if you generate a payment schedule, you'll see the payment go up at month 61

jimthompson5910 (jim_thompson5910):

from 548.09 to 626.13

jimthompson5910 (jim_thompson5910):

you're probably wondering how to get 626.13 ? Well what you do is use that table to see the remaining owed balance at month 60 (after payment is made for that month). That balance is used to recalculate the monthly payment. Instead of 360 months, it's now 300 months since 60 months have already been paid off. The value of i will increase to what was discussed above

OpenStudy (givingtree):

okay i see!!! thank you ! i did wonder.

jimthompson5910 (jim_thompson5910):

so in a sense, you're calculating the mortgage monthly payment 4 times.

OpenStudy (givingtree):

now, to filling out the chart, the # of payments column will all be 360 monthly payment column will be all $548.09 ?

jimthompson5910 (jim_thompson5910):

year1 through year5 represents 5 years (5 years)*(12 months/1 year) = 5*12 = 60 months

jimthompson5910 (jim_thompson5910):

year6 through year15 represents 10 years 10 years = 10*12 = 120 months

jimthompson5910 (jim_thompson5910):

You get the idea. Be careful about the last period

jimthompson5910 (jim_thompson5910):

also, total cost = (number of months)*(monthly payment)

OpenStudy (givingtree):

year16-25 = 9(12) ?

OpenStudy (givingtree):

no, year 16-25 would be 120 as well?

OpenStudy (givingtree):

and 26-30 would be 60 ?

OpenStudy (givingtree):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

year1 through year5 = 60 months year6 through year15 = 120 months year16 through year25 = 120 months year26 through year30 = 60 months

jimthompson5910 (jim_thompson5910):

60+120+120+60 = 360 months total

OpenStudy (givingtree):

awesome, thank you! then i just multiply the months by $548.09!

jimthompson5910 (jim_thompson5910):

for year1 through year5, yes

OpenStudy (givingtree):

oh my god im so stupid,

jimthompson5910 (jim_thompson5910):

that will give the total cost for that time period